### v2

# PWM: Policy Learning with Large World Models

Abstract
Reinforcement Learning (RL) has achieved impressive results on complex tasks but struggles in multi-task settings with different embodiments. World models offer scalability by learning a simulation of the environment, yet they often rely on inefficient gradient-free optimization methods. We introduce Policy learning with large World Models (PWM), a novel model-based RL algorithm that learns continuous control policies from large multi-task world models. By pre-training the world model on offline data and using it for first-order gradient policy learning, PWM effectively solves tasks with up to 152 action dimensions and outperforms methods using ground-truth dynamics. Additionally, PWM scales to an 80-task setting, achieving up to 27% higher rewards than existing baselines without the need for expensive online planning. Visualizations and code available at imgeorgiev.com/pwm
Figure 1: We propose PWM, a novel approach to scalable RL where we harness large multi-task world models to rapidly train model-based RL policies using efficient first-order gradients. When the world models used are regularized and avoid exploding gradients, this results in efficient policy learning in $\leq$ 10 minutes per task, while achieving higher reward in both single-task and multi-task settings.
1 Introduction
The pursuit of generalizability in machine learning has recently been propelled by the training of large models on substantial datasets (brown2020language; kirillov2023segment; bommasani2021opportunities). Such advancements have notably permeated robotics, where multi-task behavior cloning techniques have shown remarkable performance (brohan2023rt; team2023octo; goyal2023rvt; bousmalis2023robocat). Nevertheless, these approaches predominantly hinge on near-expert data and struggle with adaptability across diverse robot morphologies due to their dependence on teleoperation (brohan2023rt; team2023octo; kumar2021should).
In contrast, Reinforcement Learning (RL) offers a robust framework capable of learning from suboptimal data, addressing the aforementioned limitations. However, RL has been focused on single-task experts (mnih2013playing; schulman2017proximal; haarnoja2018soft). Recent work suggests that a potential pathway to multi-task RL is with the world models framework, where a large model learns the environment dynamics and is then combined with zeroth-order planning to solve various tasks (ebert2018visual; schrittwieser2020mastering; xu2022xtra; feng2023finetuning; hansen2023td; schubert2023generalist). Notably, TD-MPC2 (hansen2023td) performs 80 continuous control tasks using a single 317M parameter multi-task world model.
Despite advancements, zeroth-order methods often result in suboptimal solutions due to high variance (mohamed2020monte; suh2022differentiable; parmas2023model) and online planning time scales with model size, rendering them infeasible at scale (schubert2023generalist; hansen2023td). These issues are particularly noticeable in world models with billions of parameters such as GAIA-1 (hu2023gaia) and UniSim (yang2023unisim). To address these challenges, we suggest a new approach: using world models as scalable, differentiable physics simulators to enable efficient policy learning through First-order Gradients (FoG). FoGs have demonstrated better performance and stability than traditional zeroth-order methods (xu2022accelerated). We analyze world models with differentiable simulation principles and find that the utility of world models lies not in their sheer accuracy but in (1) their smoothness, (2) their optimality gap, and (3) their gradient stability over long-horizon forecasts (suh2022differentiable; howell2022dojo).
Building on these insights, we introduce Policy learning with large World Models (PWM), a novel Model-Based RL (MBRL) algorithm and framework aimed at deriving effective continuous control policies from large, multi-task world models. We utilize pre-trained TD-MPC2 world models to efficiently learn control policies with FoG in <10m per task. Our empirical evaluations on complex locomotion tasks indicate that PWM not only achieves higher reward than baselines but also outperforms methods that use ground-truth simulation dynamics. In a multi-task scenario utilizing a pre-trained 48M parameter world model, PWM achieves up to 27% higher reward than TD-MPC2 without relying on online planning.
This underscores the efficacy of PWM and supports our broader contributions:
  1. 1.
    Through pedagogical examples and ablations, we show that more accurate world models do not translate to better performing policies. Instead of pursuing world model improvements in isolation, we should aim to build world models that result in better policies.
  2. 2.
    When regularized correctly, world models enable efficient first-order optimization. We show that this results in better performing policies and faster training times in comparison to zeroth-order gradient methods.
  3. 3.
    We propose PWM, an algorithm for learning expert policies from large multi-task world models in minutes using First-order Gradients (FoG).
2 Related work
RL approaches can be classified as model-based and model-free which assume and do not assume a model respectively (arulkumaran2017deep). Most common algorithms for real-world applications such as PPO (schulman2017proximal) and SAC (haarnoja2018soft) are model-free and fall in the category of on-policy and off-policy methods respectively (arulkumaran2017deep). Both harness an actor-critic architecture where the critic learns the value and the actor optimizes the critic directly to maximize cumulative rewards (konda1999actor). Off-policy methods such as SAC typically learn a policy by taking First-order Gradients (FoG) directly from the critic. FoGs optimization typically has lower variance but is affected by discontinuities of the objective (mohamed2020monte). On-policy methods such as PPO harness zeroth-order gradients to learn policies (sutton1999policy). These gradients are not affected by discontinuities, making them effective for robotic tasks such as locomotion (rudin2022learning) but also exhibit high variance, which leads to slow optimization and suboptimal solutions (mohamed2020monte; suh2022differentiable). Differentiable simulation has been a popular framework to explore the properties of these gradient types (suh2022differentiable; howell2022dojo; xu2022accelerated).
While RL remains mostly focused on single-task policies, the larger robotics community has shifted focus to large multi-task models relying on behavior cloning (firoozi2023foundation). RT-1 (brohan2022rt) trains a large Transformer sequence model on a vast dataset to perform low-level object manipulation on a real robot. RT-2 (brohan2023rt) extends this by pre-training an even larger model on internet-scale data to also do planning and grasp affordance. Most recently, Open X (padalkar2023open) and Octo (team2023octo) showed that training these approaches on different tasks and embodiment leads to increased performance on single-task single-embodiment evaluations. Unfortunately, the benefits of large models and large datasets have not been sufficiently explored in the context of RL. DreamerV3 (hafner2023mastering) is a Model-Based Reinforcement Learning (MBRL) approach which learns a representation of its environment with encoder, decoder, dynamics and reward modules (collectively dubbed world model). It showed that scaling to larger models (400M parameters) leads to higher data efficient and episode rewards. TD-MPC2 (hansen2023td) continued this trend by employing a 317M parameter world model to learn an online planning multi-task policy capable of achieving good performance on 80 tasks. While showing impressive multi-task scalability, TD-MPC2 fails to solve all tasks it was trained on and exhibits limited scalability due to its use of online planning. With world models increasingly growing in size, we need a policy learning technique capable of scaling to billion parameter world models such as GAIA-1 (hu2023gaia) or UniSim (yang2023unisim).
3 Background
We focus on discrete-time and infinite horizon Reinforcement Learning (RL) scenarios characterized by system states ${\bm{s}}\in\mathbb{R}^{n}=\mathcal{S}$, actions ${\bm{a}}\in\mathbb{R}^{m}=\mathcal{A}$, dynamics function $f:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ and a reward function $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$. Combined, these form a Markov Decision Problem (MDP) summarized by the tuple $(\mathcal{S},\mathcal{A},f,r,\gamma)$ where $\gamma$ is the discount factor. Actions at each timestep $t$ are sampled from a stochastic policy ${\bm{a}}_{t}\sim\pi_{\theta}(\cdot|{\bm{s}}_{t})$, parameterized by ${\bm{\theta}}$. The goal of the policy is to maximize the cumulative discounted rewards:
$\max_{\bm{\theta}}J({\bm{\theta}}):=\max_{\bm{\theta}}\operatorname{\mathbb{E}}_{\begin{subarray}{c}{\bm{s}}_{1}\sim\rho(\cdot)\\ {\bm{a}}_{t}\sim\pi_{\bm{\theta}}(\cdot|{\bm{s}}_{t})\end{subarray}}\left[\sum_{t=1}^{\infty}\gamma^{t}r({\bm{s}}_{t},{\bm{a}}_{t})\right]$ (1)
where $\rho({\bm{s}}_{1})$ is the initial state distribution. Since this maximization over an infinite sum is intractable, in practice we often maximize over a value estimate. The value of a state ${\bm{s}}_{t}$ is defined as the expected reward follow the policy $\pi_{\bm{\theta}}$
$V^{\pi}_{\bm{\psi}}({\bm{s}}_{t}):=\operatorname{\mathbb{E}}_{{\bm{a}}_{h}\sim\pi_{\bm{\theta}}(\cdot|{\bm{s}}_{h})}\left[\sum_{h=t}^{\infty}\gamma^{h}r({\bm{s}}_{h},{\bm{a}}_{h})\right]$ (2)
When $V$ is approximated with a learned model with parameters ${\bm{\psi}}$ and $\pi_{\bm{\theta}}$ attempts to maximize some function of $V$, we arrive at the popular and successful actor-critic architecture (konda1999actor). Additionally, in MBRL it is common to also learn approximations of $f$ and $r$, which we denote as $F_{\bm{\phi}}$ and $R_{\bm{\phi}}$, respectively. It has also been shown to be beneficial to encode the true state ${\bm{s}}$ into a latent state ${\bm{z}}$ using a learned encoder $E_{\bm{\phi}}$ (hafner2019dream; hansen2022temporal; hansen2023td; hafner2023mastering). Putting together all of these components we can define a model-based actor-critic algorithm to consist of the tuple $(\pi_{\bm{\theta}},V_{\bm{\psi}},E_{\bm{\phi}},F_{\bm{\phi}},R_{\bm{\phi}}$) which can describe popular approaches such as Dreamer (hafner2019dream) and TD-MPC2 (hansen2023td). Notably, we make an important distinction between the types of components. We refer to $E_{\bm{\phi}}$, $F_{\bm{\phi}}$ and $R_{\bm{\phi}}$ as the world model components since they are a supervised learning problem with fixed targets. On the other hand, $\pi_{\theta}$ and $V_{\psi}$ optimize for moving targets which is fundamentally more challenging and we refer to them as the policy components.
4 Policy optimization through learned world models
This paper builds on the insight that since access to $F_{\bm{\phi}}$ and $R_{\bm{\phi}}$ is assumed through a pre-trained world-model, we have the option to optimize Eq. 1 via First-order Gradient (FoG) optimization which exhibit lower gradient variance, more optimal solutions and improved sample efficiency (mohamed2020monte). In our setting, these types of gradients are obtained by directly differentiating the expected terms of Eq. 1 as shown in Eq. 3. Note that this gradient estimator is also known as reparameterized gradient (kingma2015variational) and pathwise derivative (schulman2015gradient). While we use the explicit $\nabla^{[1]}$ notation below, we later drop it for simplicity as all gradient types in this work are first-order gradients.
$\nabla_{{\bm{\theta}}}^{[1]}J({\bm{\theta}}):=\operatorname{\mathbb{E}}_{\begin{subarray}{c}{\bm{s}}_{1}\sim\rho(\cdot)\\ {\bm{a}}_{h}\sim\pi_{\bm{\theta}}(\cdot|{\bm{s}}_{h})\end{subarray}}\left[\nabla_{\bm{\theta}}\bigg{(}\sum_{t=1}^{\infty}\gamma^{t}r({\bm{s}}_{t},{\bm{a}}_{t})\bigg{)}\right]$ (3)
As $\nabla_{{\bm{\theta}}}^{[1]}J({\bm{\theta}})$ in itself is a random variable, we need to estimate it. A popular way to do that in practice is via Monte-Carlo approximation where we are interested in two properties - bias and variance. In Sections 4.1 and LABEL:sec:learning-chaos we tackle each aspect with toy robotic control problem to build intuition. In Section LABEL:sec:method we combine our findings to propose a new algorithm.
4.1 Learning through contact
(a) Ball-wall visualization.
(b) Problem landscape.