### v3

# PWM: Policy Learning with Multi-Task World Models

Abstract
Reinforcement Learning (RL) has made significant strides in complex tasks but struggles in multi-task settings with different embodiments. World model methods offer scalability by learning a simulation of the environment but often rely on inefficient gradient-free optimization methods for policy extraction. In contrast, gradient-based methods exhibit lower variance but fail to handle discontinuities. Our work reveals that well-regularized world models can generate smoother optimization landscapes than the actual dynamics, facilitating more effective first-order optimization. We introduce Policy learning with multi-task World Models (PWM), a novel model-based RL algorithm for continuous control. Initially, the world model is pre-trained on offline data, and then policies are extracted from it using first-order optimization in less than 10 minutes per task. PWM effectively solves tasks with up to 152 action dimensions and outperforms methods that use ground-truth dynamics. Additionally, PWM scales to an 80-task setting, achieving up to 27% higher rewards than existing baselines without relying on costly online planning. Visualizations and code are available at imgeorgiev.com/pwm.
Figure 1: We propose PWM, a new method for multi-task RL that utilizes pre-trained world models to learn policies for each task. When sufficiently regularized, these world models induce smooth optimization landscapes, which allows for efficient first-order optimization. Our approach can solve tasks in <10 minutes and achieves higher rewards in both single-task and multi-task environments.
1 Introduction
The pursuit of generalizability in machine learning has recently been propelled by the training of large models on substantial datasets brown2020language; kirillov2023segment; bommasani2021opportunities. Such advancements have notably permeated robotics, where multi-task behavior cloning techniques have shown remarkable performance brohan2023rt; team2023octo; goyal2023rvt; bousmalis2023robocat. Nevertheless, these approaches predominantly hinge on near-expert data and struggle with adaptability across diverse robot morphologies due to their dependence on teleoperation brohan2023rt; team2023octo; kumar2021should.
In contrast, Reinforcement Learning (RL) offers a robust framework capable of learning from suboptimal data, addressing the aforementioned limitations. However, traditional RL has been focused on single-task experts mnih2013playing; schulman2017proximal; haarnoja2018soft. Recently, hansen2023td suggested that a potential pathway to multi-task RL is with the world models framework, where a large model learns the environment dynamics and is then combined with Zeroth-order Gradient (ZoG) methods. Despite advancements, ZoG methods struggle with sample inefficiency due to the high variance mohamed2020monte; suh2022differentiable; parmas2023model and online planning time scales with model size, rendering it infeasible at scale.
First-order Gradient (FoG) methods provide a low-variance alternative that have shown superior sample efficiency and asymptotic performance when combined with smooth differentiable simulations xu2022accelerated. However, they struggle to optimize through discontinuities suh2022differentiable; georgiev2024adaptive. In this work, we explore the tight coupling between FoG optimization and world models through the lens of differentiable simulation. Counter-intuitively, we find that for gradient-based optimization, we don’t want world models to be accurate; instead, we want them to be smooth and have a low optimality gap. This in turn enables efficient FoG optimization.
Building on these insights, we propose Policy learning with multi-task World Models (PWM), an algorithm that can learn policies from offline pre-trained world models in under <10 minutes per task. With this new-found efficiency, we also propose a new multi-task framework, where instead of training a multi-task algorithm, we only train a multi-task world model and then extract a policy for each task. This decoupling of the supervised and RL objectives results in more stable and efficient learning with higher episode rewards. Our empirical evaluations on high-dim. tasks indicate that PWM not only achieves higher reward than baselines but also outperforms methods that use ground-truth dynamics. In a multi-task scenario utilizing a pre-trained 48M parameter world model from TD-MPC2, PWM achieves up to 27% higher reward than TD-MPC2 without relying on online planning. This underscores the efficacy of PWM and supports our broader contributions:
  1. 1.
    Correlation Between World Model Smoothness and Policy Performance: Through pedagogical examples and ablations, we demonstrate that smoother, better-regularized world models significantly enhance policy performance. Notably, this results in an inverse correlation between model accuracy and policy performance.
  2. 2.
    Efficiency of First-Order Gradient (FoG) Optimization: We show that combining FoG optimization with well-regularized world models enables more efficient policy learning compared to zeroth-order methods. Furthermore, policies learned from world models asymptotically outperform those trained with ground-truth simulation dynamics, emphasizing the importance of the tight relationship between FoG optimization and world model design.
  3. 3.
    Scalable Multi-Task Algorithm: Instead of training a single multi-task policy model, we propose PWM, a framework where a multi-task world model is first pre-trained on offline data. Then per-task expert policies are extracted in <10 minutes per task, offering a clear and scalable alternative to existing methods focused on unified multi-task models.
2 Background
We focus on discrete-time and infinite-horizon Reinforcement Learning (RL) scenarios characterized by system states ${\bm{s}}\in\mathbb{R}^{n}=\mathcal{S}$, actions ${\bm{a}}\in\mathbb{R}^{m}=\mathcal{A}$, dynamics function $f:\mathcal{S}\times\mathcal{A}\rightarrow\mathcal{S}$ and a reward function $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$. Combined, these form a Markov Decision Problem (MDP) summarized by the tuple $(\mathcal{S},\mathcal{A},f,r,\gamma)$ where $\gamma$ is the discount factor. Actions at each timestep $t$ are sampled from a stochastic policy ${\bm{a}}_{t}\sim\pi_{\theta}(\cdot|{\bm{s}}_{t})$, parameterized by ${\bm{\theta}}$. The goal of the policy is to maximize the cumulative discounted rewards:
$\max_{\bm{\theta}}J({\bm{\theta}}):=\max_{\bm{\theta}}\operatorname{\mathbb{E}}_{\begin{subarray}{c}{\bm{s}}_{1}\sim\rho(\cdot)\\ {\bm{a}}_{t}\sim\pi_{\bm{\theta}}(\cdot|{\bm{s}}_{t})\end{subarray}}\left[\sum_{t=1}^{\infty}\gamma^{t}r({\bm{s}}_{t},{\bm{a}}_{t})\right]$ (1)
where $\rho({\bm{s}}_{1})$ is the initial state distribution. Since this maximization over an infinite sum is intractable, in practice we often maximize over a value estimate. The value of a state ${\bm{s}}_{t}$ is defined as the expected reward following the policy $\pi_{\bm{\theta}}$
$V^{\pi}_{\bm{\psi}}({\bm{s}}_{t}):=\operatorname{\mathbb{E}}_{{\bm{a}}_{h}\sim\pi_{\bm{\theta}}(\cdot|{\bm{s}}_{h})}\left[\sum_{h=t}^{\infty}\gamma^{h}r({\bm{s}}_{h},{\bm{a}}_{h})\right]$ (2)
When $V$ is approximated with a learned model with parameters ${\bm{\psi}}$ and $\pi_{\bm{\theta}}$ attempts to maximize some function of $V$, we arrive at the popular and successful actor-critic architecture konda1999actor. Additionally, in MBRL it is common to also learn approximations of $f$ and $r$, which we denote as $F_{\bm{\phi}}$ and $R_{\bm{\phi}}$, respectively. It has also been shown to be beneficial to encode the true state ${\bm{s}}$ into a latent state ${\bm{z}}$ using a learned encoder $E_{\bm{\phi}}$ hafner2019dream; hansen2022temporal; hansen2023td; hafner2023mastering. Putting together all of these components, we can define a model-based actor-critic algorithm to consist of the tuple $(\pi_{\bm{\theta}},V_{\bm{\psi}},E_{\bm{\phi}},F_{\bm{\phi}},R_{\bm{\phi}}$) which can describe popular approaches such as Dreamer hafner2019dream; hafner2023mastering and TD-MPC2 hansen2023td. Notably, we make an important distinction between the types of components. We refer to $E_{\bm{\phi}}$, $F_{\bm{\phi}}$ and $R_{\bm{\phi}}$ as the world model components since they are supervised learning problems with fixed targets. On the other hand, $\pi_{\theta}$ and $V_{\psi}$ optimize for moving targets, which is fundamentally more challenging, and we refer to them as the policy components.
3 Policy optimization through world models
This paper builds on the insight that since access to $F_{\bm{\phi}}$ and $R_{\bm{\phi}}$ is assumed through a pre-trained world model, we have the option to optimize Eq. 1 via First-order Gradient (FoG) optimization which exhibit lower gradient variance, more optimal solutions, and improved sample efficiency mohamed2020monte. In our setting, these types of gradients are obtained by directly differentiating the expected terms of Eq. 1 as shown in Eq. 3. Note that this gradient estimator is also known as reparameterized gradient kingma2015variational and pathwise derivative schulman2015gradient. While we use the explicit $\nabla^{[1]}$ notation below, we later drop it for simplicity as all gradient types in this work are first-order gradients.
$\nabla_{{\bm{\theta}}}^{[1]}J({\bm{\theta}}):=\operatorname{\mathbb{E}}_{\begin{subarray}{c}{\bm{s}}_{1}\sim\rho(\cdot)\\ {\bm{a}}_{h}\sim\pi_{\bm{\theta}}(\cdot|{\bm{s}}_{h})\end{subarray}}\left[\nabla_{\bm{\theta}}\bigg{(}\sum_{t=1}^{\infty}\gamma^{t}r({\bm{s}}_{t},{\bm{a}}_{t})\bigg{)}\right]$ (3)
As $\nabla_{{\bm{\theta}}}^{[1]}J({\bm{\theta}})$ in itself is a random variable, we need to estimate it. A popular way to do that in practice is via Monte-Carlo approximation, where we are interested in two properties: bias and variance. In Sections 3.1 and LABEL:sec:learning-chaos we tackle each aspect with a toy robotic control problem to build intuition. In Section LABEL:sec:method, we combine our findings to propose a new algorithm.
3.1 Learning through contact
(a) Ball-wall visualization.
(b) Problem landscape.